OpenStudy (anonymous):

Given that AE is a tangent of the circle with the center at B, BC = 3.5, and AE = 24, which is the length of EC?

OpenStudy (anonymous):

drawwwwwwwwwwwwwww it :)

OpenStudy (anonymous):

there you go

OpenStudy (anonymous):

pythagorean theorem ;)

OpenStudy (anonymous):

A. 7 B. 13 C. 25 D. 49

OpenStudy (hang254):

you have a right triangle.... the chord AC is a diameter tangents and a radius intersect at right angles so angle EAB = 90 BC = 3.5 so AC = 7 (2 radii in a diameter) EC is found by Pythagoras' theorem EC^2=AE^2+AC^2 EC^2=24^2+7^2

OpenStudy (hang254):

You can solve for the rest

Directrix (directrix):

@qudder17 Look at the attached diagram. If you know your Pythagorean triples, you can answer the question without any calculation. Otherwise, use the Pythagorean Theorem as instructed above.